# Causal Head Gating: A Framework for Interpreting Roles of Attention Heads in Transformers

### v1

@@ Abstract
We present causal head gating (CHG), a scalable method for interpreting the functional roles of attention heads in transformer models. CHG learns soft gates over heads and assigns them a causal taxonomy—facilitating, interfering, or irrelevant—based on their impact on task performance. Unlike prior approaches in mechanistic interpretability, which are hypothesis-driven and require prompt templates or target labels, CHG applies directly to any dataset using standard next-token prediction. We evaluate CHG across multiple large language models (LLMs) in the Llama 3 model family and diverse tasks, including syntax, commonsense, and mathematical reasoning, and show that CHG scores yield causal—not merely correlational—insight — validated via ablation and causal mediation analyses. We also introduce contrastive CHG, a variant that isolates sub-circuits for specific task components. Our findings reveal that LLMs contain multiple sparse, sufficient sub-circuits, that individual head roles depend on interactions with others (low modularity), and that instruction following and in-context learning rely on separable mechanisms.
@@ 1 Introduction
Large language models (LLMs) achiam2023gpt ; liu2024deepseek ; grattafiori2024llama represent the state-of-the art across a wide array of domains, exhibiting remarkable generalization and problem-solving capabilities. Yet, as these models grow in scale and complexity, they become increasingly opaque, making it more difficult to understand, predict, or control their behavior, which raises concerns about safety and misuse bommasani2021opportunities ; wei2023jailbroken ; weidinger2024holistic . This has motivated a growing body of work on interpretability, which seeks to better understand how LLMs learn and represent information, and how their responses can be shaped (tenney2019bert, ; bricken2023towards, ). Interest has focused in particular on transformer-based architectures vaswani2017attention such as GPT achiam2023gpt , LLaMA grattafiori2024llama , Gemma team2025gemma , and DeepSeek liu2024deepseek , in which the central processing blocks consist of multi-head attention followed by multi-layer perceptrons. Here, there has been considerable research on the roles of individual attention heads, which have been found to exhibit some level of human-interpretability elhage2021mathematical ; todd2023function ; yang2025emergent .
@@
Two broad categories of approaches dominate research on mechanistic interpretability in LLMs. The first uses a trained mapping from latent representations to human-interpretable concepts, such as syntactic features tenney2019bert ; tenney2019you ; hewitt2019structural or identifiable items (e.g., the Golden Gate Bridge templeton2024scaling ). The second uses causal interventions to identify portions of a single weight matrix or individual attention heads responsible for a specific behavior (lee2024mechanistic, ; voita2019analyzing, ). These approaches often focus on small portions of a model — ‘zooming in’ (olah2020zoom, ) in an effort to interpret the role of a single computational subgraph. However, in deep-learning models, computation is often distributed hinton1986learning and the role of one component is dependent on another elhage2022superposition ; fakhar2022systematic ; giallanza2024integrated , making the behavior of such complex distributed systems difficult to predict from an understanding of their parts alone (mitchell2009complexity, ).
@@
To apply a distributed perspective to mechanistic interpretability, we introduce causal head gating (CHG) which seeks to identify a parametrically weighted set of heads that contribute to a model’s execution of a given task. Given a dataset that defines a task, we fit a set of gating values for each attention head that applies a soft ablation to its output using next-token prediction, so that task-facilitating heads remain unaltered while any task-interfering heads are suppressed. Using a simple regularization procedure that further separates irrelevant heads from those that facilitate or interfere with task performance, CHG assigns meaningful scores to each attention head across an entire model according to its task contribution. We use these scores to define a taxonomy of task relevance according to how individual attention heads contribute to a model’s distributed computation of a given task, describing each head as facilitating, interfering or irrelevant. In this respect, CHG offers an exploratory complement to standard hypothesis-driven approaches to mechanistic interpretablity, assigning causal roles without relying on predefined hypotheses about what each head might be doing.
@@
Beyond its conceptual contribution, CHG also offers several practical methodological advantages over existing mechanistic interpretability tools. First, because CHG operates directly on next-token prediction, it avoids the need for externally-provided labels tenney2019bert ; tenney2019you ; hewitt2019structural ; templeton2024scaling , controlled input-output pairs tenney2019bert ; tenney2019you ; hewitt2019structural , or rigid prompt templates wang2022interpretability ; todd2023function ; yang2025emergent , which are often required for decoding and interventional approaches. Second, CHG naturally accommodates complex target outputs, including chain-of-thought reasoning wei2022chain , where the solution spans multiple intermediate steps. Finally, CHG is highly scalable: it introduces only one learnable parameter per attention head and requires no updates to the underlying model weights, so that the CHG parameters can be fitted in minutes using gradient-based optimization, even for LLMs with billions of parameters.
@@
    We use CHG to show that models contain multiple low-overlap, task-sufficient sub-circuits, suggesting head roles are not fully modular but depend on interactions with other heads.
@@
    We use CHG to show that instruction following and in-context learning rely on separable circuits at the head level, and that one can be selectively suppressed through CHG-guided head gating without disrupting the other.
@@ 2 Related Work
Representational decoders are models trained to map hidden activations to externally labeled properties tenney2019bert ; tenney2019you ; hewitt2019structural , estimating the mutual information between representations and those properties belinkov2022probing ; pimentel2020information . However, such probing results are difficult to interpret: simpler decoders may underfit and miss relevant features (false negatives), while complex decoders may overfit and learn spurious correlations (false positives) hewitt2019designing ; voita2020information , requiring complexity-accuracy tradeoffs to contextualize results voita2020information . Moreover, although decodability indicates that a property is encoded in the representation, it does not imply that the model uses that information for its task—highlighting a correlational finding rather than a causal one ravichander2020probing . Finally, representational decoders require labeled datasets, constraining their use to curated, predefined properties. For a comprehensive review of the probing framework and its limitations, see belinkov2022probing .
Sparse autoencoders can be viewed as a related approach, where the model reconstructs representations through a sparse bottleneck to reveal modular or interpretable features templeton2024scaling ; cunningham2023sparse . However, like probing classifiers, their insights remain correlational and still depend on post hoc labeling or interpretation, inheriting the same supervision bottleneck.
@@ 3.1 Applying gates to attention heads
Figure 1: (a) Schematic of a single multihead attention block with CHG-determined gating attenuation (in red). (b) Gate fitting trajectories for three heads. When fitting with $\lambda<0$ and $\lambda>0$, $G^{+}$ and $G^{-}$ both stay near 1 for facilitating heads and near 0 for interfering heads, but bifurcate to 1 and 0 respectively for irrelevant heads. (c) Gate values after fitting.
@@
Role Description $G^{+}$ $G^{-}$ Metric Ablation Effect
@@ 4.2 Distribution of causal roles
It is also worth noting that, across all tasks, 84.0% of heads are marked as facilitating or interfering (score $\geq 0.5$) in at least one seed, yet only a small fraction are consistently facilitating or interfering across all seeds (Figure 3b). In syntax and commonsense tasks, most models have fewer than 5% of heads that are always facilitating and virtually none that are always interfering (Table 2). In contrast, math reveals more rigid and consistent circuitry, with up to 38.3% of heads consistently facilitating and 1.3% consistently interfering. These patterns suggest that individual attention heads may not have modular, context-independent roles, but instead participate in a flexible ensemble of overlapping sub-circuits, in which their function depends on the configuration of others.
@@ 4.3 Comparison with causal mediation analysis
CMA, like CHG, aims to identify attention heads that facilitate task execution, though it does so in a more hypothesis-driven manner. Framed in signal detection terms, CMA and CHG are complementary. CMA exhibits high precision but relatively low sensitivity: while many facilitating heads may go undetected (false negatives), those it does identify are reliably task-relevant (few false positives) Conversely, CHG is biased toward sensitivity over precision. This suggests that that heads identified by CMA should also be identified (as showing strong facilitation) under CHG. We test this by comparing CHG to the results of two former studies using CMA, replicating their methods to identify attention heads with specific computations: heads that encode task information in function vectors todd2023function and heads that perform symbolic reasoning yang2025emergent .
@@
We apply a similar logic to symbolic reasoning tasks from yang2025emergent , where the goal is to generalize abstract identity rules such as ABA (“flow^Started^flow”) or ABB (“flow^Started^Started”). We deploy the same CMA procedure used in yang2025emergent to identify the three-stage symbolic processing mechanism that was reported: (1) symbol abstraction heads that abstract symbols (“A” or “B”) away from the actual tokens in the in-context examples; (2) symbolic induction heads that operate over the abstracted symbols to induce the symbol for the missing token in the query; (3) retrieval heads that retrieve the actual token based on the induced symbol to complete the query. To screen heads of each type, we construct prompt pairs in which either the same token is assigned to different symbols (“A” or “B”) or tokens are swapped while preserving the same rule, and patch activations at certain token positions between them. Attention heads that steer model behavior towards specific hypotheses about the three head types after patching (either converting the abstract rule or altering the actual token) are labeled as mediating. We conduct all experiments on the Llama-3.2-3B-Instruct model.
@@ 4.4 Contrastive Causal Head Gating
$\mathcal{L}_{\text{contrastive}}(G;\mathcal{M}_{\theta},\mathcal{D}_{R},\mathcal{D}_{F},\lambda)=\mathcal{L}(G;\mathcal{M}_{\theta},\mathcal{D}_{R},\lambda)-\mathcal{L}(G;\mathcal{M}_{\theta},\mathcal{D}_{F},\lambda)$ (2)
where $\mathcal{D}_{R}$ is the dataset to retain, $\mathcal{D}_{F}$ is the dataset to forget, $\lambda<0$, and $\mathcal{L}$ is the original CHG loss defined in Equation 1.
@@ 5 Discussion
In this work, we introduced Causal Head Gating (CHG), a flexible and scalable method for identifying causally relevant attention heads in large language models. CHG assigns each head a graded score for facilitation, interference, or irrelevance based on its causal effect on task performance, going beyond correlational or observational analyses. Crucially, it does so using next-token prediction alone, thereby avoiding reliance on labeled data or handcrafted prompts, making it broadly and easily applicable. Moreover, CHG requires no finetuning or auxiliary decoder model, and introduces only one parameter per head, allowing it to run in minutes even on billion-scale models.
We validated CHG by showing that its scores predict performance changes under targeted ablations, confirming that facilitation, interference, and irrelevance scores capture causal impact. Heads identified by CMA also receive high CHG facilitation scores.
Using CHG, we analyzed a range of models and tasks, revealing that attention heads form task-sufficient sub-circuits with low overlap. Interestingly, we found that head roles are not fixed: the same head may facilitate or interfere depending on which other heads are active, underscoring the distributed and context-dependent nature of computation in LLMs. Lastly, we extended CHG to a contrastive setting, showing that heads supporting instruction following and in-context learning can be selectively suppressed, indicating separable functional mechanisms.
Despite its strengths, CHG has key limitations. First, it identifies causally relevant heads but not why they matter or what they compute. Even contrastive CHG offers limited insight into underlying mechanisms, making it best suited for exploratory analysis alongside hypothesis-driven tools like CMA. Additionally, $G^{+}$ and $G^{-}$ do, albeit rarely, diverge, possibly reflecting richer head interactions. Finally, head roles can vary across runs, so multiple CHG fits may be needed.

### v2

@@ Abstract
We present causal head gating (CHG), a scalable method for interpreting the functional roles of attention heads in transformer models. CHG learns soft gates over heads and assigns them a causal taxonomy—facilitating, interfering, or irrelevant—based on their impact on task performance. Unlike prior approaches in mechanistic interpretability, which are hypothesis-driven and require prompt templates or target labels, CHG applies directly to any dataset using standard next-token prediction. We evaluate CHG across multiple large language models (LLMs) in the Llama 3 model family and diverse tasks, including syntax, commonsense, and mathematical reasoning, and show that CHG scores yield causal, not merely correlational, insight validated via ablation and causal mediation analyses. We also introduce contrastive CHG, a variant that isolates sub-circuits for specific task components. Our findings reveal that LLMs contain multiple sparse task-sufficient sub-circuits, that individual head roles depend on interactions with others (low modularity), and that instruction following and in-context learning rely on separable mechanisms.
@@ 1 Introduction
Large language models (LLMs) achiam2023gpt ; liu2024deepseek ; grattafiori2024llama represent state-of-the-art systems across a wide array of domains, exhibiting remarkable generalization and problem-solving capabilities. Yet, as these models grow in scale and complexity, they become increasingly opaque, making it more difficult to understand, predict, or control their behavior, which raises concerns about safety and misuse bommasani2021opportunities ; wei2023jailbroken ; weidinger2024holistic . This has motivated a growing body of work on interpretability, which seeks to better understand how LLMs learn and represent information, and how their responses can be shaped (tenney2019bert, ; bricken2023towards, ). Interest has focused in particular on transformer-based architectures vaswani2017attention such as GPT achiam2023gpt , LLaMA grattafiori2024llama , Gemma team2025gemma , and DeepSeek liu2024deepseek , in which the central processing blocks consist of multi-head attention followed by multi-layer perceptrons. Here, there has been considerable research on the roles of individual attention heads, which have been found to exhibit some level of human-interpretability elhage2021mathematical ; todd2023function ; yang2025emergent .
@@
Two broad categories of approaches dominate research on mechanistic interpretability in LLMs. The first uses a trained mapping from latent representations to human-interpretable concepts, such as syntactic features tenney2019bert ; tenney2019you ; hewitt2019structural or identifiable items (e.g., the Golden Gate Bridge templeton2024scaling ). The second uses causal interventions to identify portions of a single weight matrix or individual attention heads responsible for a specific behavior (lee2024mechanistic, ; voita2019analyzing, ). These approaches often focus on small portions of a model, ‘zooming in’ (olah2020zoom, ) in an effort to interpret the role of a single computational subgraph. However, in deep-learning models, computation is often distributed hinton1986learning and the role of one component is dependent on another elhage2022superposition ; fakhar2022systematic ; giallanza2024integrated , making the behavior of such complex distributed systems difficult to predict from an understanding of their parts alone (mitchell2009complexity, ).
@@
To apply a distributed perspective to mechanistic interpretability, we introduce causal head gating (CHG) which identifies a parametrically weighted set of heads that contribute to a model’s execution of a given task. Given a dataset that defines a task, we fit a set of gating values for each attention head that applies a soft ablation to its output using next-token prediction, so that task-facilitating heads remain unaltered while any task-interfering heads are suppressed. Using a simple regularization procedure that further separates irrelevant heads from those that facilitate or interfere with task performance, CHG assigns meaningful scores to each attention head across an entire model according to its task contribution. We use these scores to define a taxonomy of task relevance according to how individual attention heads contribute to a model’s distributed computation of a given task, describing each head as facilitating, interfering or irrelevant. In this respect, CHG offers an exploratory complement to standard hypothesis-driven approaches to mechanistic interpretability, assigning causal roles without relying on predefined hypotheses about what each head might be doing.
@@
Beyond its conceptual contribution, CHG also offers several practical methodological advantages over existing mechanistic interpretability tools. First, because CHG operates directly on next-token prediction, it avoids the need for externally-provided labels tenney2019bert ; tenney2019you ; hewitt2019structural ; templeton2024scaling , controlled input-output pairs tenney2019bert ; tenney2019you ; hewitt2019structural , or rigid prompt templates wang2022interpretability ; todd2023function ; yang2025emergent , which are often required for decoding and interventional approaches. Second, CHG naturally accommodates complex target outputs, including chain-of-thought reasoning wei2022chain , where the solution spans multiple intermediate steps. Finally, CHG is highly scalable: it introduces only one learnable parameter per attention head and requires no updates to the underlying model weights, so that the CHG parameters can be fitted in minutes using gradient-based optimization, even for LLMs with billions of parameters. Thus, in settings where analyzing complex dependencies between heads is important, it is feasible to fit large samples of CHG values to estimate a distribution over gating values in a bootstrap fashion.
@@
    We use CHG to show that models contain multiple task-sufficient sub-circuits with varying degrees of overlap, suggesting head roles are not fully modular but depend on interactions with other heads.
@@
    We use CHG to show that instruction following and in-context learning rely on context-dependent separable circuits at the head level, where CHG-guided gating can selectively suppress one mode without substantially disrupting the other.
The accompanying repository for this paper can be found at https://github.com/andrewnam/causal_head_gating.
@@ 2 Related Work
Representational decoders are models trained to map hidden activations to externally labeled properties tenney2019bert ; tenney2019you ; hewitt2019structural , estimating the mutual information between representations and those properties belinkov2022probing ; pimentel2020information . However, such probing results are difficult to interpret: simpler decoders may underfit and miss relevant features (false negatives), while complex decoders may overfit and learn spurious correlations (false positives) hewitt2019designing ; voita2020information , requiring complexity-accuracy tradeoffs to contextualize results voita2020information . Moreover, although decodability indicates that a property is encoded in the representation, it does not imply that the model uses that information for its task, highlighting a correlational finding rather than a causal one ravichander2020probing . Finally, representational decoders require labeled datasets, constraining their use to curated, predefined properties. For a comprehensive review of the probing framework and its limitations, see belinkov2022probing .
Sparse autoencoders (SAE) can be viewed as a related approach, where the autoencoder reconstructs representations through a sparse bottleneck to reveal modular or interpretable features templeton2024scaling ; cunningham2023sparse . However, like probing classifiers, their insights remain correlational and still depend on post hoc labeling or interpretation, inheriting the same supervision bottleneck. In contrast, CHG performs direct interventions on model components without external supervision and proposes sufficient sub-circuits to the default unablated model, thereby identifying causal links between attention heads and model behavior on a task.
@@
Most closely related to our work are differentiable masking and soft-gating approaches that learn which attention heads to retain or suppress. In de2020decisions , the authors apply sparsity gating to identify subcircuits and use the fitted parameters as weighting values in convex combinations for activation patching. Similarly, yin2024lofit learns scaling constants for each attention head, but uses the fitted values to identify heads that are most suitable for fine-tuning. Thus, while methodologically similar, our work is unique in applying the gating parameters to identify task-sufficient causal sub-circuits.
Others voita2019analyzing ; li2021differentiable have opted for hard, binary ablations using the Gumbel-softmax trick jang2016categorical ; maddison2016concrete , fitting gating probabilities rather than weighting parameters. Although these Gumbel-based approaches have been applied for causal circuit discovery in a similar spirit to our work, they suffer from a fundamental limitation that CHG does not. Specifically, while Gumbel-based gating methods also learn differentiable gates per head, they treat each head independently, effectively learning separate Gumbel–Bernoulli distributions for head inclusion. This factorized formulation models only marginal probabilities and cannot capture interdependencies between heads that jointly affect task performance. In contrast, CHG jointly optimizes all gating coefficients under the model’s loss, capturing the full range of interactions and contingencies between the attention heads. Because CHG is highly scalable, it can be fit repeatedly across random seeds or subsets, effectively sampling from the space of sub-circuits without assuming independence between heads. This enables estimation of the underlying distribution over functional head configurations while preserving the joint statistical structure that factorized gating approaches discard.
@@ 3.1 Applying gates to attention heads
Figure 1: (a) Schematic of a single multihead attention block with CHG-determined gating attenuation (in red). (b) Gate fitting trajectories for three heads on L3.2-3BI with OpenMathInstruct2. When fitting with $\lambda<0$ and $\lambda>0$, $G^{+}$ and $G^{-}$ both stay near 1 for facilitating heads and near 0 for interfering heads, but bifurcate to 1 and 0 respectively for irrelevant heads. (c) Gate values after fitting.
@@
Role Description $^{+} GG^{-}$ Metric Ablation Effect
@@ 4.2 Distribution of causal roles
It is also worth noting that, across all tasks, 84.0% of heads are marked as facilitating or interfering (score $\geq 0.5$) in at least one seed, yet only a small fraction are consistently facilitating or interfering across all seeds (Figure 3b). In syntax and commonsense tasks, most models have fewer than 5% of heads that are always facilitating and virtually none that are always interfering (Table 2). In contrast, math reveals more rigid and consistent circuitry, with up to 38.3% of heads consistently facilitating and 1.3% consistently interfering. These patterns suggest that individual attention heads may not have modular, context-independent roles, but instead participate in a flexible ensemble of overlapping sub-circuits, in which their function depends on the configuration of others merullo2024talking .
@@ 4.3 Comparison with causal mediation analysis
CMA, like CHG, aims to identify attention heads that facilitate task execution, though it does so in a more hypothesis-driven manner. Framed in signal detection terms, CMA and CHG are complementary. CMA exhibits high precision but relatively low sensitivity: while many facilitating heads may go undetected (false negatives), those it does identify are reliably task-relevant (few false positives). Conversely, CHG is biased toward sensitivity over precision. This suggests that heads identified by CMA should also be identified (as showing strong facilitation) under CHG. We test this by comparing CHG to the results of two former studies using CMA, replicating their methods to identify attention heads with specific computations: heads that encode task information in function vectors todd2023function and heads that perform symbolic reasoning yang2025emergent .
@@
We apply a similar logic to symbolic reasoning tasks from yang2025emergent , where the goal is to generalize abstract identity rules such as ABA (“flowˆStartedˆflow”) or ABB (“flowˆStartedˆStarted”). We deploy the same CMA procedure used in yang2025emergent to identify the three-stage symbolic processing mechanism that was reported: (1) symbol abstraction heads that abstract symbols (“A” or “B”) away from the actual tokens in the in-context examples; (2) symbolic induction heads that operate over the abstracted symbols to induce the symbol for the missing token in the query; (3) retrieval heads that retrieve the actual token based on the induced symbol to complete the query. To screen heads of each type, we construct prompt pairs in which either the same token is assigned to different symbols (“A” or “B”) or tokens are swapped while preserving the same rule, and patch activations at certain token positions between them. Attention heads that steer model behavior towards specific hypotheses about the three head types after patching (either converting the abstract rule or altering the actual token) are labeled as mediating. We conduct all experiments on the Llama-3.2-3B-Instruct model.
@@ 4.4 Contrastive Causal Head Gating
$\mathcal{L}(G;\mathcal{M}_{\theta},\lambda)=\sum_{(x_{R},y_{R},x_{F},y_{F})}\log P(y_{F}\mid x_{F})-\log P(y_{R}\mid x_{R})-\lambda\sum_{i,j}\sigma^{-1}(G_{l,h})$ (2)
where $\log P(y\mid x)$ denotes the log-probability of target sequence $y$ given prompt $x$ under model $\mathcal{M}_{\theta}$ with gating matrix $G$, the sum ranges over matched tuples $(x_{R},y_{R},x_{F},y_{F})$ of the retention and forget variants that differ only in task formulation, and $\lambda>0$. To stabilize the gradient, we clip the inverse-sigmoid as in Eq. 1 as well as the difference in log-probability.
@@ 5 Discussion
In this work, we introduced Causal Head Gating (CHG), a flexible and scalable method for identifying causally relevant attention heads in large language models. CHG assigns each head a graded score for facilitation, interference, or irrelevance based on its effect on task performance, going beyond correlational or observational analyses. These scores predict performance changes under targeted ablations, confirming that facilitation, interference, and irrelevance scores capture causal impact. Crucially, it does so using next-token prediction alone, thereby avoiding reliance on labeled data or handcrafted prompts, making it broadly and easily applicable. Moreover, CHG requires no finetuning or auxiliary decoder model, and introduces only one parameter per head, allowing it to run in minutes even on billion-scale models. To validate our method, we demonstrated that existing works within the mechanistic interpretability literature successfully corroborate our findings using, and that the ICL and instruction-following circuits revealed using contrastive CHG successfully generalize across tasks.
Interestingly, across the range of models and tasks we investigated, we observed that attention heads form task-sufficient sub-circuits with low overlap. Moreover, a single head may vary in its relevance across multiple runs depending on which others are active, reflecting the distributed and context-dependent nature of computation in LLMs, and in rare cases, a head may even receive low $G^{+}$ but high $G^{-}$ scores within the same run. We hypothesize that this variability reflects an interaction-dependent landscape in which causal roles shift with circuit configuration. While these complexities may appear messy, we view them as a strength of CHG, revealing the redundancy and interdependence that underlie emergent model behavior. Because CHG is highly scalable, it can be repeatedly applied to estimate distributions over gating values, providing a bootstrapped view of redundant and contingent sub-circuits with greater fidelity to the model’s underlying dependency structure.
While CHG provides a lightweight and scalable approach for exploratory analysis, requiring only a dataset and no model finetuning or supervision, it is not designed to reveal the precise computations performed by individual heads. Instead, CHG offers a complementary first-pass diagnostic tool that identifies candidate heads or sub-circuits with consistent causal influence, guiding where more granular, hypothesis-driven methods such as causal mediation or activation patching can be applied. In this way, CHG provides a practical entry point into large-scale causal interpretability, mapping functional dependencies that subsequent analyses can examine in greater detail.